A free lunch?: “We will install new HVAC equipment for you, absolutely free!! Just signa three-year contract to split 50/50 any savings the equipment generates on the use of equipment plus pay for a service contract for the equipment’s maintenance. We will takeout the loan for the equipment and installation and pay off the loan with our share of the savings. After three years, the equipment is yours! You’ll only be giving us some of the money you are paying to the power company for electricity. The service contract for the equipment is $15,000 per year, but think of your peace of mind!” Evaluate if this is a free lunch or if there is a catch somewhere based on other financial data available. What data would you need? What evaluation criteria would you choose?
1
Expert's answer
2020-12-23T07:24:27-0500

Free cheese only in the mousetrap!

The total cost of using the equipment should be compared with the amount of savings per year that the equipment brings

"15 000\times3-0.5\times15000\times3=45000-22500=22500"

Total costs are more than savings, so it's not a free lunch